# Clustering of timed sequences – Application to the analysis of care pathways

### v1

@@ 7.2 Cluster Analysis of K-means with drop-DTW
In this section, we present the results of the K-means clustering with drop-DTW metric applied on the entire cohort. Figure 4 illustrates the 5 clusters yielded. Each cluster is depicted with its barycenters and the distribution of events.
Let us first comment on the clusters based on their event distributions. Three main clusters have been identified (clusters 1, 2 and 5) and share almost the same size (about 1000 patients). The two others clusters are very small (number 3 and 4). These small clusters of patients seem to be outliers. Similarly to the previous results, the K-means identified cluster of patients with different delays between the pre-operative consultations and resection. The cluster 2 corresponds to patients having a small delay between this events and possibly two pre-operative consultations. These patients possibly have a late stage of the disease and need a quick surgery. We made this hypothesis because their survival at 6 months is significantly lower than the two others clusters. The cluster 5 and cluster 1 differs mainly in the delay between the surgery and the post-operative consultations. The Cluster 1 also holds patients without exams the day of the resection.
The same observations can be made based on the analysis of the barycenters. We remind that the barycenter is a probabilistic timed sequence. Each event is represented by a distribution of event’s types. On the figure, each bar is an event located at its time of occurrence. A bar stacks the probabilities per event’s type. We remark that most of the events have one main type. The event at time 10 for cluster 1 is a mixture of mainly two events types: post-operative consultation or post-operative multi-disciplinary consultation meeting (MCM). This possible confusion may be explained by the imprecision of the records. MCM are identified by consultation reports. If the report is not correctly tagged as an MCM, then it is simply a consultation. Then, the flexibility of the metric makes it somehow robust to possible labeling errors in the data.
We can also notice that there is a clear alignement between the barycenters events and the event distribution. This illustrates that the averaging method, based on drop-DTW provides a meaningful average of a collection of complex timed sequences. More specifically, comparing the clusters by the decay between consultations and surgery leads to the same conclusions as before.
Finally, one can remark that there are a lot of repeated events in the barycenter. Furthermore, the presence of repeated events in the barycenter can be misleading, as it could either indicate a repetition of events or alternative delays for a single event due to possible drops.
@@ 8 Conclusion
We proposed a method for clustering timed sequences inspired by existing techniques for time series. Thus, the use of timed sequences allows co-occurring events, irregular time sampling, and sparse events. All these advantages are very interesting from a clinical point of view for describing care pathways. We showed that by using drop-DTW, the algorithm became less sensitive to missing and aberrant values, which is also useful in application to real-world care pathways. Then, we tested our method on real data from the OPTISOINS study for our approach compared to TraMineR. Based on the initial results, drop-DTW seems to be clinically promising for identifying average pathways.
However, the large number of parameters and the required computation time are limitations of the method. Regarding the number of parameters, our effort focused on a guide for choosing parameter values, which the experiments tend to validate. In perspective, a more in-depth study of the effect of parameters still needs to be carried out. Additionally, we have to validate the clinical interest of average pathways by clinicians.

### v2

@@ 7.2 Cluster Analysis of K-means with drop-DTW
In this section, we present the results of the K-means clustering with drop-DTW metric applied on the entire cohort. Figure 4 illustrates the $K=5$ clusters. Each cluster is depicted with its barycenters and the distribution of events.
Let us first comment on the clusters based on their event distributions. Three main clusters have been identified (clusters 1, 2 and 3 and share almost the same size (about 1000 patients). The two other clusters are very small (numbered 4 and 5). These small clusters of patients seem to be outliers. Similarly to the previous results, the K-means identified clusters of patients with different delays between the pre-operative consultations and resection. The cluster 1 corresponds to patients having a small delay between these events and possibly two pre-operative consultations. These patients possibly have a late stage of the disease and need a quick surgery. We made this hypothesis because their survival at 6 months is significantly lower than the two others clusters. The cluster 2 and cluster 3 differ mainly in the delay between the surgery and the post-operative consultations. The cluster 3 also holds patients without exams the day of the resection.
The same observations can be made based on the analysis of the barycenters. We remind that a barycenter is a probabilistic timed sequence. Each event is represented by a distribution of event’s types. On the figure, each bar is an event located at its time of occurrence. A bar stacks the probabilities per event’s type. We remark that most of the events have one main type. The event at the 10th day for cluster 3 is a mixture of mainly two events types: post-operative consultation or post-operative multi-disciplinary consultation meeting (MCM). This possible confusion may be explained by the imprecision of the records. MCM are identified by consultation reports. If the report is not correctly tagged as an MCM, then it is simply a consultation. Then, the flexibility of the metric makes it somehow robust to possible labeling errors in the data.
We can also notice that there is a clear alignment between the barycenters events and the event distribution. This illustrates that the averaging method, based on drop-DTW provides a meaningful average of a collection of complex timed sequences. More specifically, comparing the clusters by the decay between consultations and surgery leads to the same conclusions as before.
Lastly, one can note that there are a lot of repeated events in the barycenter. Furthermore, the presence of repeated events in the barycenter can be misleading, as it could improve either indicate a repetition of events or alternative delays for a single event due to possible drops.
@@ 8 Conclusion
We proposed a method for clustering timed sequences inspired by existing techniques for time series. Thus, the use of timed sequences allows co-occurring events, irregular time sampling, and sparse events. All these advantages are very interesting from a clinical point of view for describing care pathways. We showed that by using drop-DTW, the clustering algorithm are less sensitive to missing and aberrant values when drop-DTW is used. The result is interesting for application to real-world care pathways. Then, we tested our method on real data from the OPTISOINS study for our approach compared to TraMineR. Based on the initial results, drop-DTW seems to be clinically promising for identifying average pathways.
However, the large number of parameters is a first weakness of the method. Regarding the number of parameters, our effort focused on a guide for choosing parameter values, which the experiments tend to validate. In perspective, a more in-depth study of the effect of parameters still needs to be carried out. Additionally, we have to validate the clinical interest of average pathways by clinicians.
Finally, the number of computations of a drop-DTW required during the clustering of timed sequences may be very high and may lead to long computation times for large datasets. Some strategies that could make use of distributed resources (grid computing) or fast approximate of the metric could significantly the practical usability of our method.